OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1)with a slope of 4. Answer y = 4x - 7 y = 4x + 11 y = 4x + 7 y = 4x - 11

OpenStudy (anonymous):

y+1=4(x+3) y=4x+11